ganeshie8 (ganeshie8):

Find all solutions of \[\large x^2 \equiv 11 \pmod{3167}\]

OpenStudy (perl):

x = 1417 mod 3167 x = -1417 mod 3167

OpenStudy (perl):

(this is not the most efficient approach) To solve this ,since 3167 is prime x^2 = 11 mod 3167 x^2 - 11 = 0 mod 3167 (x - sqrt(11) ) ( x + sqrt(11) = 0 mod 3167 x = sqrt(11) x = - sqrt(11) We need to find the square root of 11 modulo 3167 by squaring every number from 0 to 3167 (in fact we only need to go halfway) and reducing modulo 3167, we can see that x = 1417

ganeshie8 (ganeshie8):

Yes! those are all the solutions

OpenStudy (mathmath333):

@perl how did u find "square root of 11 mod 3167".

OpenStudy (perl):

i used this calculator http://www.numbertheory.org/php/squareroot.html

OpenStudy (perl):

enter a: 11 m : 3167

OpenStudy (mathmath333):

ok

OpenStudy (perl):

there is a better way to solve this, using number theory

ganeshie8 (ganeshie8):

wolfram is pretty good wid modular arithmetic too http://www.wolframalpha.com/input/?i=x%5E2%3D11+mod+3167

ganeshie8 (ganeshie8):

yes there is a nice simple formula..

ganeshie8 (ganeshie8):

\[x^2\equiv a \pmod{p} \iff x \equiv \pm a^{(p+1)/4} \pmod{p}\]

ganeshie8 (ganeshie8):

that works always except when \(p\) is of form \(4k+1\)

OpenStudy (perl):

oh -1417 = 1750 mod 3167

ganeshie8 (ganeshie8):

yes :) atmost two solutions are possible as the degree is \(2\)

OpenStudy (mathmath333):

i thought u used primitive roots to solve this

ganeshie8 (ganeshie8):

Exactly! that formula was derived using primitive roots

ganeshie8 (ganeshie8):

we don't need to use that formula if we know a primitive root... then we can directly take discrete logarithm both sides but we're not given a primitive root of 3167 in the problem, so...

OpenStudy (mathmath333):

this case is for prime \(p\) only or for prime and composite both \(x^2\equiv a \pmod{p} \iff x \equiv \pm a^{(p+1)/4} \pmod{p}\)

ganeshie8 (ganeshie8):

\(p\) must be a prime

ganeshie8 (ganeshie8):

see if it really works using below simple example : \[x^2 \equiv 2 \pmod{7}\]

OpenStudy (perl):

so the problem reduces to solving $$ \Large { x = \pm 11^{\frac{3167 + 1}{4}} \mod {3167}\\ x = \pm 11^{792} \mod {3167}\\ }$$

ganeshie8 (ganeshie8):

yes \(x=\pm 11^{792}\) is a legitimate answer :)

OpenStudy (perl):

is there a way to simplify that , so it is less than 3167

OpenStudy (mathmath333):

in case i know the primitive root is \(5\) then how to solve further \(\large \color{black}{\begin{align} 2\text{ind}_5x\equiv \text{ind}_511 \pmod {3166}\hspace{.33em}\\~\\ \end{align}}\)

ganeshie8 (ganeshie8):

thats a neat way to solve this! xD lets see...

ganeshie8 (ganeshie8):

solve \(\text{ind}_5x\) first

OpenStudy (mathmath333):

how

OpenStudy (mathmath333):

and how to form table for this upto 10

ganeshie8 (ganeshie8):

solving is exactly same as solving any logarithm equation

ganeshie8 (ganeshie8):

\[\large \color{black}{\begin{align} 2\text{ind}_5x\equiv \text{ind}_511 \pmod {3166}\hspace{.33em}\\~\\ \end{align}}\] is exactly same as \[\large \color{black}{\begin{align} 2\text{log}_5x\equiv \text{log}_511 \pmod {3166}\hspace{.33em}\\~\\ \end{align}}\]

ganeshie8 (ganeshie8):

how do we solve a logarithm equation ? first we need to "have" the value of \(\log_5 11\) we need to look up in logarithm table because computing it manually is a pain

ganeshie8 (ganeshie8):

basically we need the discrete logarithm table for base 5.. we can form a table but hmm... nobody cooks up logarithm tables everytime they face with a log equation... they just look up preexisting logarithm tables right

OpenStudy (mathmath333):

ok how to compute that with wolfram \(ind_5=11\)

ganeshie8 (ganeshie8):

one sec..

OpenStudy (mathmath333):

i mean \(\text{ind}_511\)

ganeshie8 (ganeshie8):

for \(\text{ind}_5 11\), weneed to solve \[11 = 5^x \pmod{316\color{red}{7}}\]

ganeshie8 (ganeshie8):

also there is a direct command for discretelogarithm : http://www.wolframalpha.com/input/?i=MultiplicativeOrder%5B5%2C+3167%2C+%7B11%7D%5D

OpenStudy (mathmath333):

ok \(2476\)

ganeshie8 (ganeshie8):

yes plug that in and solve ind_5 x

ganeshie8 (ganeshie8):

\[\large \color{black}{\begin{align} 2\text{ind}_5x\equiv \text{ind}_511 \pmod {3166}\hspace{.33em}\\~\\ \end{align}}\] \[\large \color{black}{\begin{align} 2\text{ind}_5x\equiv 2476 \pmod {3166}\hspace{.33em}\\~\\ \end{align}}\] simply divide \(2\) through out

ganeshie8 (ganeshie8):

we get \[\large \color{black}{\begin{align} \text{ind}_5x\equiv 1238 \pmod {1583}\hspace{.33em}\\~\\ \end{align}}\]

OpenStudy (mathmath333):

yes this one

ganeshie8 (ganeshie8):

yes solve x

OpenStudy (mathmath333):

\(x\equiv 5^{1238} \pmod {1583}\)

OpenStudy (mathmath333):

is that correct \(\uparrow\)

ganeshie8 (ganeshie8):

nope 1238 is the index of x in mod 3167 relative to base 5

ganeshie8 (ganeshie8):

\[x\equiv 5^{1238} \pmod{\color{Red}{3167}}\]

OpenStudy (mathmath333):

okk

OpenStudy (mathmath333):

y did u write 1583 \(\large \color{black}{\begin{align} \text{ind}_5x\equiv 1238 \pmod {\color{red}{1583}}\hspace{.33em}\\~\\ \end{align}}\)

ganeshie8 (ganeshie8):

thats right, there are no typoes

ganeshie8 (ganeshie8):

that comes from dividing 2 in : \[\large \color{black}{\begin{align} 2\text{ind}_5x\equiv 2476 \pmod {3166}\hspace{.33em}\\~\\ \end{align}}\]

OpenStudy (mathmath333):

ok

ganeshie8 (ganeshie8):

I know it is a bit confusing as we're changing modulus back and forth among 3 different numbers : 3167, 3166, 1583 it just takes some more practice to see whats going on clearly